40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Jul 19 04:44:09 2018, Annalisa, Terra, Summary, Thermal Compensation, frequency shift observed with heating! scanning_fine_2018-07-19-07-32-08_parsed.pdfscanning_fine_2018-07-19-06-57-47_parsed.pdfYscan_scanning_parsed_2am.txt.pdf
    Reply  Fri Jul 20 07:43:54 2018, Steve, Summary, SUS, ETMY  ETMY_leveling.pngETMY.png
    Reply  Sat Jul 21 01:06:49 2018, gautam, Summary, Thermal Compensation, Y arm locking 
       Reply  Sat Jul 21 14:03:19 2018, Koji, Summary, Thermal Compensation, Y arm locking 
Message ID: 14086     Entry time: Thu Jul 19 04:44:09 2018     Reply to this: 14090   14094
Author: Annalisa, Terra 
Type: Summary 
Category: Thermal Compensation 
Subject: frequency shift observed with heating! 

Annalisa, Gautum, Koji, Terra

Summary: with the reflector setup, we measured a frequency shift of the first and second order modes! First looks of shifts show 1st HOM shift ~-10 kHz, 2nd HOM shift ~-18 kHz (carrier ~4 kHz). We saw no shift with the cylinder/lenses set up.

- - - - -

Tonight we modified the cavity scan setup: the LO is provided by the Marconi which, at the same time, is also used to scan the AUX laser frequency instead of the Agilent. In order to get rid of the free running noise between Marconi and Agilent, the Marconi frequency was scanned and, point by point, the Agilent center frequency was changed accordingly. In order to speed up the process, the whole procedure was automated. The script is called AGfast.py and can be found in /users/annalisa/postVent.

One thing that helped in improving the data quality of the phase information was to set the Agilent IF bandwidth @1kHz. Not yet clear why, but it was better than having a lower bandwidth. To be further investigated.

With this setup, we made some coarse scan of the full FSR and then we "zoomed" around the main peaks in order to increase the resolution and get a more precise information about the peak frequency.

Here are the frequency ranges that we scanned:

  • carrier - central frequency: 31.73MHz; range: [31.68MHz - 31.78MHz]
  • HOM1 - central frequency: 32.88MHz; range: [32.84MHz - 32.93MHz]
  • HOM2 - central frequency: 34.03MHz; range: [33.95MHz - 34.06MHz]
  • HOM3 - central frequency: 35.18MHz; range: [35.09MHz - 35.25MHz]

We powered the heater of the lenses setup @4:55 am at 14.4V and 0.9A. Then we slightly increased the power @5:05am and the final "hot state" configuration is with heater powered at 16V and 0.9A.

With this setup we couldn't see any frequency shift

Then, at around 6:30 am we turned on the reflector setup and we measured a frequency shift of the first and second order modes. First scans show 1st HOM shift ~10 kHz, 2nd HOM shift ~18 kHz. First attachment shows carrier hot/cold, second attachment shows HOM2 hot/cold. We started to get plauged by high seismic noise. Heaters turned off at 7:45 am. Lots of scans and actual analysis to go.


gautam: about the questionable plotting -

  • 10 faint (alpha~0.3) lines are individual measurements with the reflector doing its heating. (AG4395A, 0 span, single frequency measurements plotted together).
  • charcoal line, labelled mean, is the mean of the 10 above lines.
  • bright green ("Reference") is the mean of a coarse scan (cold ETM) overlaid for comparison. 
  • "cold" - self explanatory.

My personal favourite plot is Attachment #3, which is a 5 MHz scan (cold) to identify positions of the various peaks. The power of including phase information in the analysis is clear. The second FSR on the right edge of the plot is not as prominent as the first is because the arm transmission was degrading throughout the measurement. For future measurements, we should consider locking the IMC length to the arm cavity - this would eliminate such alignment drifts, and maybe also make the PLL control signal RMS smaller. 

Attachment 1: scanning_fine_2018-07-19-07-32-08_parsed.pdf  50 kB  Uploaded Thu Jul 19 08:47:19 2018  | Hide | Hide all
scanning_fine_2018-07-19-07-32-08_parsed.pdf
Attachment 2: scanning_fine_2018-07-19-06-57-47_parsed.pdf  51 kB  Uploaded Thu Jul 19 08:47:50 2018  | Hide | Hide all
scanning_fine_2018-07-19-06-57-47_parsed.pdf
Attachment 3: Yscan_scanning_parsed_2am.txt.pdf  110 kB  Uploaded Fri Jul 20 12:39:36 2018  | Hide | Hide all
Yscan_scanning_parsed_2am.txt.pdf
ELOG V3.1.3-